OpenStudy (anonymous):

Need help please, No direct answer please explain how to solve. Will medal Mason worked 21 hours at the bicycle shop at an hourly rate of $9.50. How much Medicare tax will be deducted from his paycheck this week? (Medicare tax is a fixed tax of 1.45% of a person’s gross pay.) A. $199.50 B. $8.38 C. $2.89 D. $196.61

OpenStudy (jdoe0001):

well Mason will be making 9.50 per hour he worked for 21 hours so 9.50 + 9.50 + .... 21 times or 21 * 9.50 is how much Mason made, in gross pay, that is, gross with no deductibles now, the taxman cometh the taxman bites off that 1.45% off the gross pay so, how much is 1.45% of that? well \(\bf {\color{brown}{ 1.45}}\%\ of\ (21\cdot 9.50) \implies \cfrac{{\color{brown}{ 1.45}}}{100}\cdot (21\cdot 9.50)\implies {\color{brown}{ 1.45}}\div 100\cdot (21\cdot 9.50)\)

OpenStudy (anonymous):

it will be c right

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

thanks

OpenStudy (jdoe0001):

yw